For this problem, carry at least four digits after the decimal in your calculations. Answers may vary slightly due to rounding.

A. In a marketing survey, a random sample of 986 supermarket shoppers revealed that 272 always stock up on an item when they find that item at a real bargain price.

Let p represent the proportion of all supermarket shoppers who always stock up on an item when they find a real bargain. Find a point estimate for p. (Round your answer to four decimal places.)

Find a 95% confidence interval for p. (Round your answers to three decimal places.)

lower limit    
upper limit    

B. In a survey of 1000 large corporations, 250 said that, given a choice between a job candidate who smokes and an equally qualified nonsmoker, the nonsmoker would get the job.

Let p represent the proportion of all corporations preferring a nonsmoking candidate. Find a point estimate for p. (Round your answer to four decimal places.)

Find a 0.95 confidence interval for p. (Round your answers to three decimal places.)

lower limit    
upper limit    
Answer #1

a) Point estimate = x/n = 272/986 = 0.2759

For 95% confidence, z = 1.96

Hence,

95% confidence interval will be:

Lower limit = 0.2759 - 0.0279 = 0.248

Upper limit = 0.2759 + 0.0279 = 0.304

b) Point estimate = x/n = 250/1000 = 0.2500

For 95% confidence, z = 1.96

Hence,

95% confidence interval will be:

Lower limit = 0.250 - 0.0268 = 0.223

Upper limit = 0.250 + 0.0268 = 0.277